2012 January 15 UT LBC/MODS checkout

Afternoon was spent recovering from a power outage this morning.

Doug Officer and I loaded the INAF set of masks.

First half of night cloudy. Dome closed.

MODS:

  • MODS computers were brought back up
    • M1.RC did not boot without intervention. Message was MOS checksum error - Defaults loaded and at the bottom press F1 to continue DEL to enter setup. After talking to Rick, pressed F1 to continue. Booted fine after that.
  • Closed Dome: taking images of MOS masks (VFLAT 3.0 and 1 sec is good for this), and some calibrations through the MOS masks.

LBC:

  • LBC came up OK
    • Noticed horizontal banding on some biases taken early on. Low-level.
    • Edited OBs:
      • Fixed exin FOCUS OBs which, for Blue side, had incorrect center point (-0.2mm instead of -0.02mm)
      • Edited SUPERFOC OBs to take 7 exposures, 10-sec each, with a step size of 0.06mm. The longer baseline and exposure times should give better results, as was seen earlier this month for RB_rVsuperfoc7. These all are called RB_xxsuperfoc7.ob and are in /home/LBTO/Calib_OBs/SUPERFOC.
    • Running an OB with dome closed but exposures long enough to trigger trackers.
      • RB_Tech is not working... RB_Science is working.
      • First image of preset has PA different from rest by about 2 degrees, for both Blue and Red channels.
Sandbox> hsel /newdata/lbcr.*.06*.fits[2] $I,telra,teldec,dithoffx,dithoffy,pa_pnt,rotangle,filter
/newdata/lbcr.20120115.061110.fits[2]   06:46:31.999    +44:51:16.98    0       0       2.64361 -162.42627      z-SLOAN
/newdata/lbcr.20120115.061323.fits[2]   06:46:31.997    +44:51:16.98    0       0       0.86176 -164.57959      i-SLOAN
/newdata/lbcr.20120115.061458.fits[2]   06:46:31.997    +44:51:16.98    0       0       0.90453 -166.09131      Y-FAN

Sandbox> hsel /newdata/lbcb.*.06*.fits[2] $I,telra,teldec,dithoffx,dithoffy,pa_pnt,rotangle,filter
/newdata/lbcb.20120115.061127.fits[2]   06:46:31.999    +44:51:16.99    0       0       2.65768 -162.42350      g-SLOAN
/newdata/lbcb.20120115.061338.fits[2]   06:46:31.998    +44:51:16.98    0       0       0.88331 -164.49332      r-SLOAN
/newdata/lbcb.20120115.061519.fits[2]   06:46:31.997    +44:51:16.97    0       0       0.93173 -166.13688      SDT_Uspe

  • /home/lbccontrol was 96% full. Moved all log files from 2011 to /images/lbccontrol_oldlogs after gzipping the ones not already compressed. /home/lbccontrol now down to 87%.
  • Bias levels from a set of 10 biases taken around 2am:
    • Blue:
      • 487-493
      • 478
      • 483-485
      • 487-488
    • Red:
      • 486-491
      • 482-485
      • 483-485
      • 483-487
06:30 UT May be clearing somewhat. Thinking about opening.

09:30 UT Doesn't look like we'll be opening. Humidity has risen, and clouds have thickened. We'll switch to MODS to run a closed-dome preset. But... already there is a problem with the LDG rotator. It is in a fault. Geno saw and resolved this earlier by reseatnig the fiber among other things. We tried these but to no avail.
  • On the phone with Dan Cox - Geno and he resolved the issue. It was a SERDES error. Geno noted also that the unit in the AUX control room (to which the fiber plugs in) seems to be sensitive to heat and so the cabinet doors, front and back, are being left open as a rule now.

11:00 UT Configured for MODS.
  • Geno sent a preset from IRAF. I sent an active preset from a MODS script. These worked except for the "failed to find guide star in image".
  • Also we imaged all of the MOS masks.

11:25 UT Humidity has risen to 87% and weather station reports "rain". In fact it is starting to snow. We're slewing to zenith and calling it a night (at 12:15 UT).
  • executed modsSleep.pro
  • turned off LBC (except for HK).

-- OlgaKuhn - 15 Jan 2012
Topic revision: r3 - 15 Jan 2012, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback